4t^2-7t=0
Determine the nature of solutions
choose from
2 real solutions
1 real solution
2 imaginary solutions

Answer by ewatrrr(24785)

 
Hi
4t^2-7t=0
t(4t-7)=0
t = 0
t = 7/4 |2 real solutions